OpenStudy (anonymous):

Solve. a + 7 = 1 a = A. 8 B. –8 C. –6 D. 6

sammixboo (sammixboo):

C

OpenStudy (uri):

a + 7 = 1 a=1-7 a=-6 ans.

OpenStudy (anonymous):

thx

sammixboo (sammixboo):

a + 7 = 1 a + 7 - 7 = 1 - 7 a = -6 (That is they way I learned how to do the problem)

OpenStudy (anonymous):

can u

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

adding